# 2Zhongguancun Laboratory, Beijing, China

### v1

@@ 2.1 LLM Ensemble Taxonomy
(b) Ensemble during inference. As the most granular form of ensemble among the three broad categories, this type of approach encompasses: (a1) token-level ensemble methods, integrate the token-level outputs of multiple models at the finest granularity of decoding; (a2) span-level ensemble methods, perform ensemble at the level of a sequence fragment (e.g., a span of four words); (a3) process-level ensemble methods, select the optimal reasoning process step-by-step within the reasoning chain for a given complex reasoning task. Note that for these ensemble-during-inference methods, the aggregated text segments will be concatenated with the preceding text and fed back into the models.
@@ 3 Methodology
inference (Sec. 3.2),text width=7.6em [(b1) Token-level ensemble,text width=9.6em [Aggregation-based: 1) Vocabulary alignment: GaC Yu et al. (2024) / DeePEn Huang et al. (2024) /
EVA Xu et al. (2024) / 2) Others: UniTe Yao et al. (2024) / PackLLM Mavromatis et al. (2024) / CDS Jin et al. (2024) ] [Selection-based: CITER Zheng et al. (2025) ] ] [(b2) Span-level ensemble, text width=9.6em [Cool-Fusion Liu et al. (2024b) / SweetSpan Xu et al. (2025) / SpecFuse Lv et al. (2024b) ] ] [(b3) Process-level ensemble, text width=10.1em [LE-MCTS Park et al. (2024) ] ] ] [(c) Ensemble-after-
inference (Sec. 3.3),text width=7.6em [(c1) Non-cascade,text width=6.4em [Selection-based: Agent-Forest Li et al. (2024) / Smoothie Guha et al. (2024) / MoRE Si et al. (2023)] [Selection-then-regeneration: LLM-Blender Jiang et al. (2023) / URG Lv et al. (2024a) / LLM-TOPLA Tekin et al. (2024) ] ] [(c2) Cascade, text width=5.4em [Unsupervised: EcoAssistant Zhang et al. (2023) / Yue et al. (2024) / Model Cascading Varshney and Baral (2022) /
@@ 3.1 Ensemble Before Inference
Unlike learning-based routers, Eagle Zhao et al. (2024) is a training-free router, which employs the ELO ranking to update the scores of LLMs based on pairwise comparison results. For each comparison, the observed scores and the expected scores are calculated to determine the score changes of the LLMs, thereby evaluating the relative skill levels of the LLMs and providing the highest quality responses to users within the given budget. PickLLM Sikeridis et al. (2024) adopts reinforcement learning to dynamically route queries to the most appropriate LLMs. By designing a weighted reward function, it comprehensively considers multiple factors such as cost, query time, and response accuracy. Blending Lu et al. (2024c) introduces a simple yet effective conversational AI method, which randomly selects different responses of multiple chat AIs, breaking the traditional model of relying on a single chat model. It effectively combines the advantages of multiple small models and significantly improves the performance of chat AI while maintaining a low inference cost.
@@ 3.2 Ensemble During Inference
Furthermore, there are several alternative methods that do not prioritize dictionary alignment as their core objective: UniTe Yao et al. (2024) and PackLLM Mavromatis et al. (2024) directly adopt existing alignment techniques, and CDS Jin et al. (2024) employs homogeneous models in their context without the need for alignment. Specifically: (i) Yao et al. Yao et al. (2024) propose UniTe, which focuses solely on the TOP-K portion of each model’s output distribution and employs the aforementioned union dictionary-based alignment strategy Yu et al. (2024) for performing averaging aggregation, avoiding the need for full vocabulary alignment and reducing computational overhead. (ii) Mavromatis et al. Mavromatis et al. (2024) present a weighted averaging approach, where the weights were determined by the perplexity of each model at that specific time step, serving as an indicator of whether the current input aligns with the respective model’s expertise. Here we present the formal definition of perplexity (considering that it is a crucial concept also employed in several other related methods Liu et al. (2024b); Xu et al. (2025)). Given a tokenized input prompt or sequence $S=(x_{0},x_{1},\ldots,x_{t})$,
@@
where the perplexity value is determined by the exponentiated average negative log-likelihood of $S$, and $\log p_{k}(x_{i}\mid x_{<i})$ denotes the log-likelihood of $i$-token given the preceding tokens $x_{<i}$ according to model $M_{k}$. Further, a relatively complex variant of this approach is derived from a perplexity-based greedy optimization problem. (iii) CDS Jin et al. (2024) ensembles the outputs of multiple homogeneous models (derived from a single ancestor model after pre-training or fine-tuning/alignment) at key decoding points (determined by a critical token classifier), enabling the simultaneous utilization of their complementary advantages to enhance the factuality of the generated content.
@@
(c1) Agent-Forest Li et al. (2024) Unsupervised Similarity-based selection $>2$ Scaling property Cls./Generation [Link]
@@ 3.3 Ensemble After Inference
We review the non-cascade and cascade methods in the following subsections, and provide a summary analysis in Table 3.
@@
As shown in Figure 2 and Table 3, existing non-cascade methods fall into two categories: 1) selection-based, which is dedicated to selecting a single response from multiple candidates Li et al. (2024); Guha et al. (2024); Si et al. (2023), and 2) selection-then-regeneration, which involves initially selecting a subset of candidate responses and subsequently feeding this refined subset into a generative model for regeneration to obtain the final output Jiang et al. (2023); Tekin et al. (2024); Lv et al. (2024a).
@@
First, as unsupervised methods, Agent-Forest Li et al. (2024) and Smoothie Guha et al. (2024) essentially leverage the similarity between model responses for selection and adhere to a majority voting (MV) principle: selecting the response that exhibits the highest degree of similarity to the others as the final answer. Specifically, (i) Li et al. Li et al. (2024) obtain multiple responses by repeatedly querying a single model (equivalent to utilizing outputs from multiple homogeneous models) and employ the MV ensemble, discovering a scaling property associated with ensembling more responses; they use BLEU scores and occurrence frequencies to quantify the similarities between responses for generation and classification tasks, respectively. (ii) Guha et al.  Guha et al. (2024) input multiple $<$query, response$>$ pairs into SentenceBERT Reimers (2019), leveraging the Euclidean distance between the resulting low-dimensional encoded vectors as a measure of similarity; and their method is theoretically supported by a constructed graphical model.
@@
As one most straightforward method, Zhang et al. Zhang et al. (2023) propose a cascade method, EcoAssistant, that relies on user judgment rather than machine algorithms to determine whether the obtained result are satisfactory and whether the cascading inference should be terminated.
@@
Specifically, (i) Jitkrittum et al. Jitkrittum et al. (2024) use the training data to learn a cascade method called post-hoc deferral; and the core idea is that, in contrast to aforementioned class-uncertainty cascade methods Varshney and Baral (2022); Ramírez et al. (2023); Dekoninck et al. (2024) which solely focus on the class uncertainty of the current model’s output, it also takes into account the estimated uncertainty of the next, stronger model’s output. (ii) Further, Chen et al. Chen et al. (2023a) and Gupta et al. Gupta et al. (2024) use the training data to learn a scoring function that produces a confidence score regarding the current model’s output, enabling cascading judgments. The key difference lies in that Chen et al. Chen et al. (2023a) utilize input features contributed by $<$query, response$>$ pairs during the scoring function learning process; whereas Gupta et al. Gupta et al. (2024), aiming to better integrate the class-uncertainty idea in general generation tasks, employ “Quantile” features constructed from token-level class uncertainty information within the response in the binary classification-based scoring function learning. (iii) Both Aggarwal et al. Aggarwal et al. (2023) and Hu et al. Hu et al. (2024a) adopt the aforementioned idea of “introducing routing in the cascading process” Dekoninck et al. (2024) and train an MDP for decision-making. Aggarwal et al. Aggarwal et al. (2023) directly determine both “whether to terminate the cascade and which stronger model to route to” within the MDP, whereas Hu et al. Hu et al. (2024a) only decide “which stronger model to route to” in the MDP and additionally rely on an answer-consistency strategy for cascading decisions.

### v2

@@ 2.1 LLM Ensemble Taxonomy
(b) Ensemble during inference. As the most granular form of ensemble among the three broad categories, this type of approach encompasses: (b1) token-level ensemble methods, integrate the token-level outputs of multiple models at the finest granularity of decoding; (b2) span-level ensemble methods, perform ensemble at the level of a sequence fragment (e.g., a span of four words); (b3) process-level ensemble methods, select the optimal reasoning process step-by-step within the reasoning chain for a given complex reasoning task. Note that for these ensemble-during-inference methods, the aggregated text segments will be concatenated with the preceding text and fed back into the models.
@@ 3 Methodology
inference (Sec. 3.2),text width=7.6em [(b1) Token-level ensemble,text width=9.6em [Aggregation-based: 1) Vocabulary alignment: GaC Yu et al. (2024) / DeePEn Huang et al. (2024) / EVA Xu et al. (2024)
/ 2) Others: UniTe Yao et al. (2024) / PackLLM Mavromatis et al. (2024) / CDS Jin et al. (2024) / Li et al. (2024b) ] [Selection-based: CITER Zheng et al. (2025) ] ] [(b2) Span-level ensemble, text width=9.6em [Cool-Fusion Liu et al. (2024b) / SweetSpan Xu et al. (2025) / SpecFuse Lv et al. (2024b) ] ] [(b3) Process-level ensemble, text width=10.1em [LE-MCTS Park et al. (2024) ] ] ] [(c) Ensemble-after-
inference (Sec. 3.3),text width=7.6em [(c1) Non-cascade,text width=6.4em [Selection-based: Agent-Forest Li et al. (2024a) / Smoothie Guha et al. (2024) / MoRE Si et al. (2023)] [Selection-then-regeneration: LLM-Blender Jiang et al. (2023) / URG Lv et al. (2024a) / LLM-TOPLA Tekin et al. (2024) ] ] [(c2) Cascade, text width=5.4em [Unsupervised: EcoAssistant Zhang et al. (2023) / Yue et al. (2024) / Model Cascading Varshney and Baral (2022) /
@@ 3.1 Ensemble Before Inference
Unlike learning-based routers, Eagle Zhao et al. (2024) is a training-free router, which employs the ELO ranking to update the scores of LLMs based on pairwise comparison results. For each comparison, the observed scores and the expected scores are calculated to determine the score changes of the LLMs, thereby evaluating the relative skill levels of the LLMs and providing the highest quality responses to users within the given budget. PickLLM Sikeridis et al. (2024) adopts reinforcement learning to dynamically route queries to the most appropriate LLMs. By designing a weighted reward function, it comprehensively considers multiple factors such as cost, query time, and response accuracy. Blending Lu et al. (2024c) introduces a simple yet effective conversational AI method that randomly selects responses from multiple small chat AIs, deviating from the traditional reliance on a single chat model and enhancing performance while maintaining low inference costs.
@@
Li et al. (2024b) Token Weighted-averaging agg. $=2$ Reduce negative issues Generation -
@@ 3.2 Ensemble During Inference
Furthermore, there are several alternative methods that do not prioritize dictionary alignment as their core objective: UniTe Yao et al. (2024) and PackLLM Mavromatis et al. (2024) directly adopt existing alignment techniques, CDS Jin et al. (2024) and Li et al.  Li et al. (2024b) employ homogeneous models in their context without the need for alignment. Specifically: (i) Yao et al. Yao et al. (2024) propose UniTe, which focuses solely on the TOP-K portion of each model’s output distribution and employs the aforementioned union dictionary-based alignment strategy Yu et al. (2024) for performing averaging aggregation, avoiding the need for full vocabulary alignment and reducing computational overhead. (ii) Mavromatis et al. Mavromatis et al. (2024) present a weighted averaging approach, where the weights were determined by the perplexity of each model at that specific time step, serving as an indicator of whether the current input aligns with the respective model’s expertise. Here we present the formal definition of perplexity (considering that it is a crucial concept also employed in other related methods Liu et al. (2024b); Xu et al. (2025)). Given a tokenized input prompt or sequence $S=(x_{0},x_{1},\ldots,x_{t})$,
@@
where $\log p_{k}(x_{i}\mid x_{<i})$ is the log-likelihood of $i$-token given the preceding tokens $x_{<i}$ according to model $M_{k}$. Further, a relatively complex variant of this approach is derived from a perplexity-based greedy optimization problem. (iii) CDS Jin et al. (2024) ensembles the outputs of multiple homogeneous models (derived from a single ancestor model after pre-training or fine-tuning/alignment) at key decoding points (determined by a critical token classifier), enabling the utilization of their complementary advantages to enhance the factuality of the generation. (iv) Focusing on a different goal, Li et al.  Li et al. (2024b) reduce negative issues during the deployment of LLMs, such as copyright infringement and data poisoning, by integrating the token-level outputs of a benign small model with a large model using weighted aggregation.
@@
(c1) Agent-Forest Li et al. (2024a) Unsupervised Similarity-based selection $>2$ Scaling property Cls./Generation [Link]
@@ 3.3 Ensemble After Inference
We review the non-cascade and cascade methods in the following, and provide a summary analysis in Table 3.
@@
As shown in Figure 2 and Table 3, existing non-cascade methods fall into two categories: 1) selection-based, which is dedicated to selecting a single response from multiple candidates Li et al. (2024a); Guha et al. (2024); Si et al. (2023), and 2) selection-then-regeneration, which involves initially selecting a subset of candidate responses and subsequently feeding this refined subset into a generative model for regeneration to obtain the final output Jiang et al. (2023); Tekin et al. (2024); Lv et al. (2024a).
@@
First, as unsupervised methods, Agent-Forest Li et al. (2024a) and Smoothie Guha et al. (2024) essentially leverage the similarity between model responses for selection and adhere to a majority voting (MV) principle: selecting the response that exhibits the highest degree of similarity to the others as the final answer. Specifically, (i) Li et al. Li et al. (2024a) obtain multiple responses by repeatedly querying a single model (equivalent to utilizing outputs from multiple homogeneous models) and employ the MV ensemble, discovering a scaling property associated with ensembling more responses; they use BLEU scores and occurrence frequencies to quantify the similarities between responses for generation and classification tasks, respectively. (ii) Guha et al.  Guha et al. (2024) input multiple $<$query, response$>$ pairs into SentenceBERT Reimers (2019), leveraging the Euclidean distance between the resulting low-dimensional encoded vectors as a measure of similarity; and their method is theoretically supported by a constructed graphical model.
@@
As one most straightforward method, Zhang et al. Zhang et al. (2023) propose a cascade method, EcoAssistant, that utilizes user judgment rather than machine algorithms to determine whether the obtained result are satisfactory and whether the cascading inference should be terminated.
@@
  • 2
    §: As a specific example and as introduced in Section 3.2.1, Li et al.  Li et al. (2024b) primarily aim to minimize the negative issues of large models during deployment.
@@
Specifically, (i) Jitkrittum et al. Jitkrittum et al. (2024) use the training data to learn a cascade method called post-hoc deferral; and the core idea is that, in contrast to aforementioned class-uncertainty cascade methods Varshney and Baral (2022); Ramírez et al. (2023); Dekoninck et al. (2024) which solely focus on the class uncertainty of the current model’s output, it also takes into account the estimated uncertainty of the next, stronger model’s output. (ii) Further, Chen et al. Chen et al. (2023a) and Gupta et al. Gupta et al. (2024) use the training data to learn a scoring function that produces a confidence score regarding the current model’s output, enabling cascading judgments. The key difference lies in that Chen et al. Chen et al. (2023a) utilize input features contributed by $<$query, response$>$ pairs during the scoring function learning process; whereas Gupta et al. Gupta et al. (2024), aiming to better integrate the class-uncertainty idea in general generation tasks, employ “Quantile” features constructed from token-level class uncertainty information within the response in the binary classification-based scoring function learning. (iii) Both Aggarwal et al. Aggarwal et al. (2023) and Hu et al. Hu et al. (2024a) adopt the aforementioned idea of “introducing routing in the cascading process” Dekoninck et al. (2024) and train an MDP for decision-making. Aggarwal et al. Aggarwal et al. (2023) directly determine both “whether to terminate the cascade and which stronger model to route to” within the MDP, whereas Hu et al. Hu et al. (2024a) decide “which stronger model to route to” in the MDP and additionally implement an answer-consistency strategy for cascading decisions.